OpenStudy (anonymous):

Find the product of (4x + 3y)^2

OpenStudy (anonymous):

i thought we help you on this one

OpenStudy (anonymous):

no it was diffrent

OpenStudy (anonymous):

oh lol ok hold on

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

(a + b)² = a² + 2 ab + b² <=> (4x + 3y)² = (4x)² + 2 (4x) (3y) + (3y)² =

OpenStudy (anonymous):

16x2 + 24xy + 9y2

OpenStudy (anonymous):

@blackprincesssen

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

r u sure 100%?!?!?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

cuz my mom showed me this 2 days ago so ik its right

OpenStudy (anonymous):

okay thnx :)

OpenStudy (anonymous):

np